OpenStudy (anonymous):

Factor the following expression 18u^3v^2x^6−12u^7v^8

OpenStudy (anonymous):

Oops - it should be 3u^9v^4w^3+15u^2v^5

myininaya (myininaya):

Hi. @Juliette2120 do you know the greatest common factor of 18 and 12?

myininaya (myininaya):

oh sorry didn't see the new problem.

myininaya (myininaya):

Do you know the greatest common factor of 3 and 15?

myininaya (myininaya):

if I write 15 as 3(5) and you compare 3 to 3(5) can you tell me the factors that 3 and 3(5) have in common?

OpenStudy (anonymous):

3 is the greatest common factor

myininaya (myininaya):

yes now let's look at the u's of each term \[3u^9v^4w^3+15u^2v^5\\ 3(u^9v^4w^3+5u^2v^5)\] so this is what we have now but we can factor out more than just what we factored out

myininaya (myininaya):

now looking at the u part... we have u^9 and u^2

myininaya (myininaya):

\[u^9=u^2 \cdot u^7\] what factors does \[u^2 \cdot u^7 \text{ and } u^2 \text{ have in common }\]

myininaya (myininaya):

let me know if you are having trouble?

OpenStudy (anonymous):

Sorry! OS won't seem to let me post my replies!

myininaya (myininaya):

I guess it let's you reply after refreshing the page ? Or do you have to completely leave and come back to post a reply? I'm on google chrome and haven o problems. Maybe you can try that internet browser.

OpenStudy (anonymous):

I think I was able to figure this out. Thank you very much for your help! Just the simplification sent me in the right direction.

myininaya (myininaya):

No problem.

myininaya (myininaya):

But we still needed to factor out a few more things.